# Mingyu Yang, Mehdi Rezagholizadeh, Guihong Li , Vikram Appia, Emad Barsoum Advanced Micro Devices, Inc. (AMD)

### v1

@@ 4.3.4 Scaling the Teacher
Model and Setting Teacher Size ARC ARE HS MM OB PI RA WG Avg.
@@ 5 Conclusion
In this work, we addressed the growing need for efficient LLMs by proposing a practical and scalable framework for composing hybrid models from existing pre-trained Transformers. Motivated by the cost and environmental impact of retraining large models for downstream use, we introduced Zebra-Llama, a family of 1B, 3B, and 8B hybrid models built using SSMs and MLA layers. We developed an effective initialization scheme and a post-training knowledge transfer pipeline that enabled these hybrid models to inherit capabilities from larger teacher models with minimal additional training. Our approach significantly reduced memory requirements while preserving or exceeding the accuracy of strong baselines, demonstrating the viability of post-training hybridization as an efficient alternative to full retraining.
@@
Our work opens several directions for future research. One limitation is the current focus on a single family of base models; in future work, we plan to explore hybridization strategies across diverse LLM architectures, including multimodal models and Mixture-of-Experts (MoE) frameworks. We also aim to scale our training pipeline to support larger models and extend our method to reasoning-intensive architectures. Another limitation lies in our reliance on strong teacher models for knowledge transfer. In some cases, a sufficiently large teacher may not be available for a given base model. Even when such teachers exist, using a separate teacher model for distillation adds to the memory and resource requirements during training. Addressing these issues—such as through teacher-free distillation or efficient self-distillation—remains an important avenue for future work.
@@ A.1 Structured MLA Initialization
[…]
@@ Appendix B SMART Layer Selection Algorithms
1:Sensitivity scores $\{s_{1},s_{2},\ldots,s_{L}\}$, number of MLA layers $N$
@@
4:Divide $L$ layers into $N$ equal partitions
5:$L_{1}^{\text{MLA}}\leftarrow$ highest-sensitivity layer in first partition of layers $\left\{i,i\in\left[1,\left\lfloor\frac{L}{N}\right\rfloor\right]\right\}$
6:$L_{N}^{\text{MLA}}\leftarrow$ highest-sensitivity layer in last partition of layers $\left\{i,i\in\left[L-\left\lfloor\frac{L}{N}\right\rfloor+1,L\right]\right\}$
@@
11:Let $R\leftarrow$ list of candidate intermediate layers in $[r_{\text{start}},r_{\text{end}}]$
@@
18:for all combinations of $N-2$ layers from $R$ do
@@
Table 7: Actually sensitivity score in Figure 3.
@@
[…]
@@ Appendix C Comparison with Pre-training Methods
In Table 8, we include more comparisons of our Zebra-Llama with other hybrid models based on pre-training instead of distillation. It can be observed that our method could achieve competitive performance as the pre-trained baselines 214×500×214$\times$-500$\times$214 × - 500 × fewer training tokens, demonstrating our advantage in the training efficiency. For instance, Zebra-Llama (16MLA–16M2) matches or exceeds the accuracy of Mamba-2-Hybrid (8.66B, 3.5T tokens) and SAMBA (1.7B, 3.5T tokens), while using only 11B tokens and smaller model sizes. This highlights the efficiency and practicality of our post-training hybrid composition strategy, as it achieves high performance with a fraction of the training budget required by scratch-trained models.
Model and Setting Tokens Size ARC ARE HS MM OB PI RA WG
@@ Appendix D More Inference Evaluations
Figure 7: Inference throughput vs. output sequence length of various 3B-size models. We measure the throughput under batch size 1.
In this section, we provide more inference evaluations on our Zebra-Llama 3B and 8B models. In Figure 7, we include the throughput evaluations for 3B models following the same setup with Figure 4 where the prompt length and batch size are fixed to 1. Similarly, our Zebra-Llama exhibits much better inference throughput than the hybrid GQA-Mamba model MambaInLlama. With 8k generation length, our Zebra-Llama is 2.26×2.26$\times$2.26 × faster than MambaInLlama with 14 MLA/GQA layers and 6.33×6.$3$ faster than MambaInLlama with 6 MLA/GQA layers.
In Figure 8 and 9, we include the inference evaluations for 8B models iunder a setting where the prompt length is fixed at 1, the output length at 8192, and the batch size gradually increases from 8 to 256. All measurements are recorded on a single AMD MI300 GPU with 192GB of memory, using the generation function from the mamba_ssm library. Our Zebra-Llama models demonstrate significant throughput gains compared to Llama, pure MLA models, and hybrid GQA-Mamba models. Although pure Mamba models achieve faster inference speeds than Zebra-Llama, they tend to exhibit poorer overall performance, as shown in Tables 1 and 2. Additionally, the KV cache compression in our Zebra-Llama models translates to notable memory savings, as illustrated in Figure 9. Models incorporating GQA layers experience a sharp increase in memory usage as the batch size grows. In contrast, MLA-based approaches, such as X-EcoMLA and our Zebra-Llama models, demonstrate superior memory efficiency, exhibiting only a marginal increase in memory consumption compared to the pure Mamba baseline.

### v2

@@ 4.3.4 Scaling the Teacher
Model and Setting Teacher Size ARC ARE HS $M$ OB PI RA WG Avg.
@@
4.4 Extension to Qwen models
[…]
Model and Setting Teacher Tokens KV Size ARC ARE HS $M$ OB PI RA WG Avg.
Qwen2.5-0.5B-Inst - - 100% 33.11 59.05 52.26 45.86 34.2 70.62 32.06 56.27 47.93
Zebra-Qwen-0.5B, 4MLA-20M2 1.5B 7B 6.25% 38.74 66.92 50.83 38.43 37.2 69.91 32.34 55.09 48.68
Qwen2.5-1.5B-Inst - - 100% 47.01 75.84 68.24 60.13 41 76.01 37.8 62.67 58.59
Zebra-Qwen-1.5B, 14MLA-14M2 1.5B 7B 6.25% 48.63 75.17 67.64 53.87 41.6 75.73 38.66 64.01 58.16
Table 6: Zero-shot evaluation on the LM Harness Eval benchmark for Zebra-Qwen models.
4.5 Extended Long-Context Evaluation
Model KV% 4K 8K 16K
MambaInLlama-1B 50% 38.75 21.55 3.88
Zebra-Llama-1B 3.91% 35.75 24.80 13.37
MambaInLlama-3B 50% 41.71 22.62 0.88
Zebra-Llama-3B 4.69% 58.69 38.24 9.97
Table 7: RULER benchmark results at 4K, 8K, and 16K context lengths.
For the long-context evaluation of our models, a few factors should be considered. First, assessing long-context performance requires models that have been trained on datasets with extended sequence lengths. However, due to efficiency and resource constraints, our current models were trained with a maximum sequence length of 2048 tokens. Such a training configuration is generally not optimal for long-context benchmarks like RULER [13]. Despite this limitation, we conducted RULER benchmark evaluations at 4K, 8K, and 16K context lengths and the results are summarized in Table 7. We observe that, at the 3B scale, Zebra-Llama surpasses MambaInLlama across all context lengths, while at the 1B scale it performs on par at 4K and exceeds MambaInLlama at 8K and 16K. Importantly, these gains are achieved with a substantially smaller memory footprint—Zebra-Llama uses only $\sim$5% of the KV cache, whereas MambaInLlama requires nearly ten times more KV memory.
As future work, we aim to extend our models to support longer context lengths and reasoning.
@@ 5 Conclusion
In this work, we addressed the growing need for efficient LLMs by proposing a practical and scalable framework for composing hybrid models from existing pre-trained Transformers. Motivated by the cost and environmental impact of retraining large models for downstream use, we introduced Zebra-Llama, a family of 1B, 3B, and 8B hybrid models built using SSMs and MLA layers. We developed an effective initialization scheme and a post-training knowledge transfer pipeline that enabled these models to inherit capabilities from larger teacher models with minimal additional training. Our approach significantly reduced memory while preserving the accuracy of strong baselines.
@@
Our work establishes several promising directions for future research. A primary next step is to move beyond a single family of base models and explore hybridization strategies across diverse LLM architectures, particularly incorporating modular frameworks like Mixture-of-Experts (MoE). Scaling our training pipeline to larger models and extending the approach to reasoning-intensive architectures remain critical frontiers. Furthermore, addressing the reliance on strong teacher models is essential. In constrained scenarios, investigating alternative knowledge transfer strategies—such as teacher-free or efficient self-distillation—will be key to understanding how hybrid architectures can effectively learn with reduced supervision.
@@ A.1 Structured MLA Initialization
where the function $\text{reshape}(.)$ is used to integrate the last two dimensions of the specified tensor.
@@ Appendix B SMART Layer Selection Algorithms
1:Sensitivity scores $\{s_{1},s_{2},\ldots,s_{L}\}$, number of MLA layers NN
@@
4:Divide $L$ layers into NN equal partitions
[…]
@@
11:Let $R\leftarrow$ list of candidate intermediate layers in $r_{\text{start}},r_{\text{end}}]$
@@
18:for all combinations of N2N-2 layers from $R$ do
@@
Table 10: Concrete sensitivity score in Figure 3.
@@
Terminal Preservation: $L^{MLA}_{1}=0$, $L^{MLA}_{N}=14$ Near-Uniform Intermediate Distribution: Define intermediate layer range: $r_{start}=1$, $r_{end}=13$, $R=\{1,2,\ldots,13\}$. Compute allowable gap bounds: $T=9$, $g_{min}=1$, $g_{max}=2$. Enumerate Valid Configurations: $\mathcal{C}=\{$ $\{2,5,8,11\}$, $\{3,5,8,11\}$, $\{3,6,8,11\}$, $\{3,6,9,11\}$, $\{3,6,9,12\}\}$. Maximal Sensitivity Scores: Calculate total sensitivity score: * $s_{2}+s_{5}+s_{8}+s_{11}=1315.5$, * $s_{3}+s_{5}+s_{8}+s_{11}=1185.8$, * $s_{3}+s_{6}+s_{8}+s_{11}=1068.4$, * $s_{3}+s_{6}+s_{9}+s_{11}=950.86$, * $s_{3}+s_{6}+s_{9}+s_{12}=890.65$. Layers with maximal score: $C^{*}=\{2,5,8,11\}$. Return $\{0,2,5,8,11,14\}$.
@@ Appendix C Comparison with Pre-training Methods
In Table 11, we include more comparisons of our Zebra-Llama with other hybrid models based on pre-training instead of distillation. It can be observed that our method could achieve competitive performance as the pre-trained baselines $214\times-500\times$ fewer training tokens, demonstrating our advantage in the training efficiency. For instance, Zebra-Llama (16MLA–16M2) matches or exceeds the accuracy of Mamba-2-Hybrid (8.66B, 3.5T tokens) and SAMBA (1.7B, 3.5T tokens), while using only 11B tokens and smaller model sizes. This highlights the efficiency and practicality of our post-training hybrid composition strategy, as it achieves high performance with a fraction of the training budget required by scratch-trained models.
Model and Setting Tokens Size ARC ARE HS $M$ OB PI RA WG
@@ Appendix D More Inference Evaluations
In Figure 8 and 9, we include more inference evaluations for 8B models iunder a setting where the prompt length is fixed at 1, the output length at 8192, and the batch size gradually increases from 32 to 1024. All measurements are recorded on a single AMD MI300 GPU with 192GB of memory with Hipgraph. Our Zebra-Llama models demonstrate significant throughput gains compared to Llama, X-EcoMLA (pure MLA), and MambaInLlama (hybrid GQA-Mamba). Additionally, the KV cache compression in our Zebra-Llama models translates to notable memory savings, as illustrated in Figure 9. Models incorporating GQA layers experience a sharp increase in memory usage as the batch size grows. In contrast, MLA-based approaches, such as X-EcoMLA and our Zebra-Llama models, demonstrate superior memory efficiency.